OpenStudy (anonymous):

A new drug store is being built at the coordinates (6, 10). Is there a street parallel to Oak Street (the entrance street) that will go to the drug store? If not, what is the equation needed to put this street on the map? PLEASE HELP

OpenStudy (anonymous):

1 we can't do without the map 2. the line y=(-5/2)x+6 has a slope of -5/2. Any parallel line will also have a slope of -5/2. So, it will have the equation y=(-5/2)x+b. Since you have a point on the line (-4,-6), plug in those for x and y and find out what b is: -6=(-5/2)*(-4)+b -6=10+b b=-16 the equation is y=(-5/2)x-16 3. The slope of the given line is 2. To find the slope of the perpendicular line, take the negative reciprocal, or -1/2. this gives you y=(-1/2)x+b. Since you have a point (4,-2), plug in those values for x and y and find b -2=(-1/2)*4+b -2=-2+b b=0 y=(-1/2)x+0 or y=(-1/2)x _/